# OWSM-CTC: An Open Encoder-Only Speech Foundation Model for Speech Recognition, Translation, and Language Identification

### v1

@@ 4.4 Speech translation
We evaluate the ST performance using the CoVoST-2 (Wang et al., 2020a) test sets. Again, we perform greedy decoding and calculate BLEU scores using lowercase without punctuation. For X-to-En translation, we follow OWSM v3.1 (Peng et al., 2024) to report results of directions where the training data size is over 100 hours. For the other low-resource directions, both OWSM v3.1 and our OWSM-CTC do not work in general. For En-to-X translation, we report results in all 15 directions. We calculate the speed-up based on the average decoding time on the NIVIDA A40 GPU.
Table 5 shows the X-to-En results. Notably, our encoder-only OWSM-CTC consistently outperforms the encoder-decoder OWSM v3.1 by a large margin. The average BLEU score is improved from 20.1 to 25.2 (25% relatively). We also achieve 3.35x speed-up for inference.
Table 6 presents En-to-X results. OpenAI Whisper does not support these directions. Similarly, our OWSM-CTC achieves superior performance than OWSM v3.1 in 12 out of 15 translation directions. The average BLEU is improved from 13.3 to 15.0 (13% relatively), and the inference speed-up is 4.20 times.
@@
Table 7: Long-form ASR results on the TEDLIUM (Hernandez et al., 2018) test set which consists of 11 audio recordings ranging from 6 to 27 minutes. Bold: the best result. Underlined: our OWSM-CTC outperforms OWSM v3.1 medium.
@@ 4.5 Long-form speech recognition
For long-form ASR, a model takes as input an unsegmented audio recording of arbitrary length and generates the entire transcription without explicit voice activity detection. Whisper and encoder-decoder OWSM can predict start and end timestamps of each utterance within a fixed-length segment. Those timestamps are used to shift the recognition window for chunk-wise long-form ASR. However, this chunk-wise recognition is a sequential process because the location of the next chunk depends on the predicted timestamp in the current chunk. By contrast, our OWSM-CTC performs chunk-wise recognition in a fully parallel manner. We first split the entire audio into overlapped chunks of 30s, where the overlapped region serves as the left and right context. [^2] We follow this tutorial for long-form ASR with CTC: https://github.com/NVIDIA/NeMo/blob/main/tutorials/asr/Streaming_ASR.ipynb We then perform CTC greedy decoding on batched chunks. The batch size is 32 on a single NVIDIA A40 GPU (48GB). Table 7 shows the WER and speed-up with different context lengths. Our OWSM-CTC achieves lower WERs than the encoder-decoder OWSM v3.1, while being approximately 20 times faster due to the batched parallel decoding. OWSM-CTC is also robust to different context lengths. These observations indicate that CTC-based non-autoregressive models perform very well for long-form ASR, which is consistent with prior findings (Koluguri et al., 2023).
Dataset Previous text as prompt? WER % ($\downarrow$)
TEDLIUM dev No 4.9
Yes 4.1
Table 8: Using the previous sentence (groundtruth) as a text prompt improves the ASR WER of OWSM-CTC. The optional prompt encoder is defined in Figure 2 and Section 3.3.
@@ 4.6 Effect of text prompt
As described in Figure 2 and Section 3.3, OWSM-CTC can take an additional text prompt as input which might change the output. During training, either a special token <na> or the previous sentence in the same audio is used as the prompt according to a probability of 0.5, which follows the setup of Whisper and OWSM. To verify that OWSM-CTC can utilize information from the prompt when necessary, we perform greedy decoding on the TEDLIUM dev set, where the previous sentence of each utterance is available. As shown in Table 8, using the previous sentence as the text prompt reduces the WER from 4.9% to 4.1%. Appendix C provides an example where the previous sentence also affects the output text style.
@@ 4.7 Robustness
To investigate the robustness, we first consider random noise as input. Table 9 shows the ASR outputs generated by three models. Encoder-decoder models including Whisper and OWSM v3.1 tend to generate some text that looks meaningful, while our OWSM-CTC only generates some punctuation marks without actual meaning. Note that punctuation marks are typically removed before ASR scoring, so our error rate will be zero.
Another typical issue for autoregressive decoding is that the generation might fall into an infinite loop of a few characters or words until reaching the maximum output length. Table 16 in Appendix D presents two examples from ASR and ST, respectively. Our non-autoregressive model is more robust in such cases.
@@ 5 Conclusion
We propose OWSM-CTC, a novel encoder-only speech foundation model built upon 180k hours of public audio data and open-source toolkits. OWSM-CTC employs multi-task self-conditioned CTC for multilingual ASR, any-to-any ST, and LID. We conduct extensive experiments to compare OWSM-CTC with the encoder-decoder OWSM models trained on the same data. We find that OWSM-CTC achieves competitive performance on ASR and superior performance on ST for both X-to-En (25% relative improvement) and En-to-X (13% relative improvement), while being more robust and 3 to 4 times faster at inference time. Additionally, OWSM-CTC improves the long-form ASR WER with 20 times faster inference due to the batched parallel decoding. OWSM-CTC also outperforms the baselines on LID. To promote open research on large speech models, we will publicly release our codebase, pre-trained model weights and training logs.
6 Acknowledgements
@@ Appendix A Details of Experimental Setups
Model Unlabelled English ASR Other ASR ST Languages Vocabulary Size
@@
 large-v3 4M hours 1M hours of labelled in total 100 52k
@@
Table 10: Details of training data. Our training data is prepared using the scripts released by OWSM v3.1 (Peng et al., 2024).
@@
Table 12: Training hyperparameters. We mostly follow the training config of OWSM v3.1 medium (Peng et al., 2024). As described in Section 3.2, we employ self-conditioned CTC at four intermediate layers.
A.1 Training data
Table 10 summarizes the training data statistics. We prepare the training data mixture using the scripts publicly released by OWSM v3.1 (Peng et al., 2024). This ensures fair comparison between our OWSM-CTC and the previously released encoder-decoder OWSM models.
A.2 Model architectures
A.3 Training hyperparameters
Table 12 presents the training hyperparameters of OWSM v3.1 and our OWSM-CTC. Again, we follow the previous OWSM v3.1 (Peng et al., 2024) for fair comparison, except that we adopt self-conditioned CTC (Nozaki and Komatsu, 2021) at four intermediate layers (see Section 3.2).
Appendix B Small-Scale Ablation Studies
@@
Table 13: Comparison of different downsampling strategies on MuST-C v2 En-De. The other configurations such as batch size are kept the same. Using 4x downsampling achieves the best ASR and ST results, while using 8x downsampling significantly reduces the GPU memory usage, which enables a larger batch size per GPU. We employ 8x downsampling in our large-scale OWSM-CTC to reduce training cost.
@@
Before the large-scale training using the entire 180k hours of audio data, we also conducted preliminary experiments on MuST-C v2 En-De (Cattoni et al., 2021) to investigate the effect of the CNN downsampling rate and the choice of the task for intermediate CTC layers. Specifically, we train 24-layer E-Branchformer-CTC models on the combined ASR and ST data from MuST-C v2 En-De. The input is always English audio, but the output can be the English ASR transcript or its German translation depending on the task specifier (see Figure 2).
@@ B.1 Effect of downsampling strategies
Table 13 compares different downsampling strategies while the other configurations are kept the same. The attention is implemented with FlashAttention (Dao et al., 2022). Self-conditioned CTC is applied at three intermediate layers: 6, 12, and 18. The first two CTC layers always perform ASR, while the others are task-dependent. The results show that using 8x downsampling in the CNN module leads to a slight degradation on WER and BLEU but reduces the GPU memory usage by a half. We thus decide to employ 8x downsampling in our large-scale OWSM-CTC, enabling a doubled batch size per GPU. As mentioned in Appendix A.2, with this strategy, we observe a similar training speed compared to the encoder-decoder OWSM model.
@@ B.2 Choice of the CTC task
As discussed in Section 3.2, the intermediate CTC layers can be configured to perform a specific task like ASR or multiple tasks depending on the input task token. Table 14 compares different choices at a small scale using MuST-C v2 En-De. If all CTC layers are task-dependent (i.e., multi-tasking), the model cannot converge when trained from scratch. As more layers are used for ASR only, the ASR WER is improved, but the ST BLEU is slightly decreased. A good trade-off is to use the first half for ASR only and the second half for multi-tasking. Therefore, in our large-scale OWSM-CTC with 27 layers, we configure the 6th, 12th and 15th layers to perform ASR only and the other two CTC layers (i.e., 21st and 27th layers) to be multi-tasking. This design also mimics the conventional cascaded system for ST.
Appendix C Effect of text prompt
Input audio content
Previous sentence
ASR w/o previous
ASR w/ previous
future ’s over here wind sun a new energy grid new investments to create high paying jobs repower america it ’s time to get real there is an old african proverb that says if you want to go quickly go alone if you want to go far go together we need to go far quickly thank you very much
with one hundred percent clean electricity within ten years a plan to put america back to work make us more secure and help stop global warming finally a solution that ’s big enough to solve our problems repower america find out more this is the last one it ’s about repowering america one of the fastest ways to cut our dependence on old dirty fuels that are killing our planet
Future’s over here. Wind, sun. A new energy grid. New investments to create high-pan jobs. Repower America. It’s time to get real. There’s an old African proverb that says, "If you want to go quickly, go alone. if you want to go far, go together." We need to go far quickly. Thank you very much. (Applause)
future ’s over here wind sun a new energy grid new investments to create high pan jobsrepower america it ’s time to get real there ’s an old african proverb that says if you want to go quickly go alone if you want to go far go together we need to go far quickly thank you very much
@@ Appendix C Effect of text prompt
Table 15 is an example from the TEDLIUM dev set, which shows that the text prompt can potentially change the output style. When there is no prompt, the ASR output of OWSM-CTC is in true case with punctuation, and the apostrophes are combined with the previous words. However, when the previous sentence is used as a prompt, the style of the ASR hypothesis becomes more similar to that of the prompt. Specifically, the text is now in lower case without punctuation marks, and the apostrophes are separate from previous words. This style is more consistent with the groundtruth transcript.
@@
Although the above example looks promising for biasing the model’s output towards certain directions, we note that this is not guaranteed to work in a zero-shot manner. We have also tried zero-shot contextual biasing, where we provide a few biasing words in the prompt (e.g., person names), but we find that the model may not be able to generate the correct word in many cases. This is mainly because the model is not really trained to perform this type of tasks - we just provide the previous sentence (according to some probability) as the prompt during training, which might not be useful at all; thus, our non-autoregressive model can simply ignore it in most cases. A more practical way to utilize this feature is to fine-tune our pre-trained model using some carefully designed data for contextual biasing. We will explore this in the future.
Groundtruth reference OWSM v3.1 output OWSM-CTC output (ours)
in search of the mythical treasure your grandfather is supposed to have secreted there he laughed and the girl instinctively shuddered with a newborn distrust there was no mirth in the sound
in search of the mythical treasure your grandfather is supposed to have secreted there ha ha ha ha ha ha ha ha ha ha ha ha ha ha ha ha ha ha …
in search of the mythical treasure your grandfather is supposed to have secreted there he laughed and the girl instinctively shuddered with a new-born distrust there was no mirth in the sound
and with her they began a national tour that took them all around the country
they take a national gira which leads to rerererererererererererererere …
with learn a national tour that leads them to run the entire country
Table 16: Autoregressive decoding sometimes gets trapped in an infinite loop for both ASR (row 1, MLS en) and ST (row 2, CoVoST-2 es-en). Our OWSM-CTC is more robust.
Appendix D Robustness

### v2

@@ 4.4 Speech translation
We evaluate ST on CoVoST-2 test sets (Wang et al., 2020a). By default, we perform greedy decoding and calculate BLEU scores in true case with punctuation. [^4] Results in lowercase without punctuation can be found in Appendix C, which are consistent with previous OWSM work (Peng et al., 2024). For X-to-En translation, we follow OWSM v3.1 (Peng et al., 2024) to report results of directions where the training data size is over 100 hours. For the other low-resource directions, both OWSM v3.1 and our OWSM-CTC do not work in general. For En-to-X translation, we report all 15 directions. We calculate the speed-up based on the average decoding time on an NIVIDA A40 GPU.
Table 5 shows the X-to-En results. Notably, our encoder-only OWSM-CTC consistently outperforms the encoder-decoder OWSM v3.1 by a large margin. The average BLEU score is improved from 20.2 to 25.1 (24% relatively). We also achieve 3.35x speed-up for inference.
Table 6 presents En-to-X results. Whisper does not support these directions. Our OWSM-CTC achieves superior performance than OWSM v3.1 in 12 of 15 translation directions and most of them are statistically significant. The average BLEU is improved from 14.1 to 16.0 (13% relatively), and the inference speed-up is 4.20 times.
@@
Table 7: Long-form ASR results on the TEDLIUM (Hernandez et al., 2018) test set which consists of 11 audio recordings ranging from 6 to 27 minutes. Bold: the best result. Underlined: OWSM-CTC outperforms OWSM v3.1 medium.
@@ 4.5 Long-form speech recognition
For long-form ASR, a model takes as input an unsegmented audio recording of arbitrary length and generates the entire transcription without explicit voice activity detection. Whisper and encoder-decoder OWSM can predict start and end timestamps of each utterance within a fixed-length segment. Those timestamps are used to shift the recognition window for chunk-wise long-form ASR. However, this chunk-wise recognition is a sequential process because the location of the next chunk depends on the predicted timestamp in the current chunk. [^5] The decoding process might be parallelized if token-level timestamps are available. However, it remains an open problem to derive accurate token-level timestamps from an attention-based encoder-decoder model without extra training. By contrast, our OWSM-CTC performs chunk-wise recognition in a fully parallel manner. We first split the entire audio into overlapped chunks of 30s, where the overlapped region serves as the left and right context. [^6] We follow this tutorial for long-form ASR with CTC: https://github.com/NVIDIA/NeMo/blob/main/tutorials/asr/Streaming_ASR.ipynb We then perform CTC greedy decoding on batched chunks. The batch size is 32 on a single NVIDIA A40 GPU (48GB). Table 7 shows the WER and speed-up with different context lengths. Our OWSM-CTC achieves lower WERs than the encoder-decoder OWSM v3.1, while being approximately 20 times faster due to the batched parallel decoding. OWSM-CTC is also robust to different context lengths. These observations indicate that CTC-based non-autoregressive models perform very well for long-form ASR, which is consistent with prior findings (Koluguri et al., 2023).
GigaSpeech LS-clean LS-other SWBD TEDLIUM AISHELL
w/o prev 11.80 2.42 5.22 16.92 4.95 6.37
w/ prev 11.23 2.38 5.10 16.70 4.55 6.25
p-value <0.001 0.19 0.007 <0.001 <0.001 <0.001
Table 8: Using the previous sentence as a text prompt improves the ASR WER/CER of OWSM-CTC.
@@ 4.6 Effect of text prompt
As described in Figure 2 and Section 3.3, OWSM-CTC can take an additional text prompt as input which might change the output. During training, either a special token <na> or the previous sentence in the same audio is used as the prompt according to a probability of 0.5, which follows the setup of Whisper and OWSM. To verify that OWSM-CTC can utilize information from the prompt when necessary, we perform greedy decoding on several test sets with the previous sentence in the dataset as a prompt. As shown in Table 8, using the previous sentence reduces the error rates. The p-values are computed using the Matched Pair Sentence Segment method. [^7] https://github.com/usnistgov/SCTK Appendix D provides an example where the previous sentence also affects the output text style.
@@ 4.7 Robustness
To investigate the robustness, we first consider random noise as input. Table 9 shows the ASR outputs generated by three models. Encoder-decoder models, including Whisper and OWSM v3.1, tend to generate some texts that look meaningful, while our OWSM-CTC generates fewer tokens, which are mostly punctuation marks that do not actually have meaning.
Another typical issue of autoregressive decoding is that the generation might fall into repetitions of a few characters or words. Table 19 in Appendix E presents two examples from ASR and ST, respectively. Our non-autoregressive model is more robust in such cases. To quantitatively measure this type of error, we consider a hypothesis as a failure if it contains any character-level $\theta$-gram ($\theta=1,2,\dots,\theta_{\text{max}}$) that consecutively occurs for at least $\delta$ times. Table 10 shows the number of failures in all ST test sets with different thresholds. We can see that the encoder-decoder OWSM v3.1 medium fails many times even with beam search, while our OWSM-CTC has almost no failures.
$\theta_{\text{max}}$ $\delta$ Model #Failures ($\downarrow$)
10 5 OWSM v3.1 2448
OWSM v3.1 (beam 5) 630
OWSM-CTC (ours) 3
20 5 OWSM v3.1 2537
OWSM v3.1 (beam 5) 672
OWSM-CTC (ours) 3
20 6 OWSM v3.1 1985
OWSM v3.1 (beam 5) 453
OWSM-CTC (ours) 1
Table 10: Comparison of the number of decoding failures in all ST test sets. There are 286k samples in total.
@@ 5 Conclusion
We propose OWSM-CTC, a novel encoder-only speech foundation model built upon 180k hours of public audio data and open-source toolkits. OWSM-CTC employs multi-task self-conditioned CTC for multilingual ASR, any-to-any ST, and LID. We conduct extensive experiments to compare OWSM-CTC with the encoder-decoder OWSM models trained on the same data. We find that OWSM-CTC achieves competitive performance on ASR and superior performance on ST for both X-to-En (24% relative improvement) and En-to-X (13% relative improvement), while being more robust and 3 to 4 times faster at inference time. Additionally, OWSM-CTC improves the long-form ASR WER with 20 times faster inference due to the batched parallel decoding. OWSM-CTC also outperforms the baselines on LID. To promote open research on large speech models, we will publicly release our code, pre-trained model weights and training logs.
Limitations
Although OWSM-CTC reduces the training cost by 22% compared to OWSM v3.1, it still requires nearly 20k GPU hours, which is nontrivial. OWSM-CTC can generate incorrect ASR or ST outputs due to limited training data in certain languages. Care should be taken when using our model for low-resource ASR or ST. Besides, we have only evaluated our model with greedy decoding as it has the fastest inference speed. The non-autoregressive model sometimes makes mistakes in spelling or grammar due to a lack of language models.
Broader Impacts and Ethics
Our OWSM-CTC is a novel encoder-only speech foundation model built upon public datasets and open-source toolkits. Compared to other popular choices, it achieves very strong performance and efficiency. We adhere to the ACL ethics policy and there is no violation of privacy in our experiments. We plan to publicly release all scripts, pre-trained models, and training logs, which can promote transparency and open science. We believe this will benefit the entire speech research community and it can make the latest speech technology available to a broader range of people all over the world.
Acknowledgements
@@ Appendix A Details of Experimental Setups
Model Unlabeled English ASR Other ASR ST Languages Vocabulary Size
@@
   large-v3 4M hours 1M hours of labeled in total 100 52k
@@
Table 11: Details of training data. Our data is prepared using the scripts released by OWSM v3.1 (Peng et al., 2024).
@@
Table 13: Training hyperparameters. We mostly follow the training setups of OWSM v3.1 medium (Peng et al., 2024). As described in Section 3.2, we employ self-conditioned CTC at four intermediate layers.
@@
Table 14: Comparison of different downsampling strategies on MuST-C v2 En-De. The other configurations, such as batch size, are kept the same. Using 4x downsampling achieves the best ASR and ST results, while using 8x downsampling significantly reduces the GPU memory usage, which enables a larger batch size per GPU. We employ 8x downsampling in our large-scale OWSM-CTC to reduce training costs.
@@
Src Lang. de es fr ca Average ($\uparrow$) Speed-up ($\uparrow$)
data size 4.3 6.7 4.5 0.2
Whisper (encoder-decoder) (Radford et al., 2023)
base 11.4 19.2 13.1 9.7 13.4 1.84x
small 25.0 32.8 26.4 21.7 26.5 1.54x
medium 33.6 39.7 34.4 29.2 34.2 0.84x
data size 0.2 0.1 0.3 0.1
OWSM v3.1 (encoder-decoder) (Peng et al., 2024)
base 7.3 10.0 11.1 9.0 9.4 2.78x
medium 17.1 22.3 22.7 18.4 20.1 1.00x
OWSM-CTC (ours)
medium 21.1 28.2 27.7 23.7 25.2 3.35x
Table 16: BLEU ($\uparrow$) of X-to-En ST on CoVoST-2 using lowercase without punctuation. Data sizes are in thousand hours. Bold: the best result. Underlined: our OWSM-CTC outperforms OWSM v3.1 medium.
data size 14.0 0.4 13.7 0.8 0.4 0.4 0.9 0.9 0.4 0.4 0.4 0.4 1.0 0.4 0.4
OWSM v3.1 (encoder-decoder) (Peng et al., 2024)
base 14.6 7.7 14.5 3.0 1.8 1.0 1.2 1.6 8.1 1.3 0.7 0.0 8.7 5.1 4.5 4.9 2.39x
OWSM-CTC (ours)
Table 17: BLEU ($\uparrow$) of En-to-X ST on CoVoST-2 using lowercase without punctuation. Data sizes are in thousand hours. Bold: the best result. Underlined: our OWSM-CTC outperforms OWSM v3.1 medium. Note that Whisper does not support En-to-X translation.
Input audio content Previous sentence ASR w/o previous ASR w/ previous
future ’s over here wind sun a new energy grid new investments to create high paying jobs repower america it ’s time to get real there is an old african proverb that says if you want to go quickly go alone if you want to go far go together we need to go far quickly thank you very much with one hundred percent clean electricity within ten years a plan to put america back to work make us more secure and help stop global warming finally a solution that ’s big enough to solve our problems repower america find out more this is the last one it ’s about repowering america one of the fastest ways to cut our dependence on old dirty fuels that are killing our planet Future’s over here. Wind, sun. A new energy grid. New investments to create high-pan jobs. Repower America. It’s time to get real. There’s an old African proverb that says, "If you want to go quickly, go alone. if you want to go far, go together." We need to go far quickly. Thank you very much. (Applause) future ’s over here wind sun a new energy grid new investments to create high pan jobsrepower america it ’s time to get real there ’s an old african proverb that says if you want to go quickly go alone if you want to go far go together we need to go far quickly thank you very much
Groundtruth reference OWSM v3.1 output OWSM-CTC output (ours)
in search of the mythical treasure your grandfather is supposed to have secreted there he laughed and the girl instinctively shuddered with a newborn distrust there was no mirth in the sound in search of the mythical treasure your grandfather is supposed to have secreted there ha ha ha ha ha ha ha ha ha ha ha ha ha ha ha ha ha ha … in search of the mythical treasure your grandfather is supposed to have secreted there he laughed and the girl instinctively shuddered with a new-born distrust there was no mirth in the sound
and with her they began a national tour that took them all around the country they take a national gira which leads to rerererererererererererererere … with learn a national tour that leads them to run the entire country
Table 19: Autoregressive decoding sometimes gets trapped in a loop in both ASR (row 1, MLS En) and ST (row 2, CoVoST-2 Es-En). Our OWSM-CTC is more robust.
A.1 Training data
Table 11 summarizes the training data statistics. We prepare the training data mixture using the scripts publicly released by OWSM v3.1 (Peng et al., 2024). This ensures a fair comparison between our OWSM-CTC and the previously released encoder-decoder OWSM models.
A.2 Model architectures
A.3 Training hyperparameters
Table 13 presents the training hyperparameters of OWSM v3.1 and our OWSM-CTC. Again, we follow the previous OWSM v3.1 (Peng et al., 2024) for a fair comparison, except that we adopt self-conditioned CTC (Nozaki and Komatsu, 2021) at four intermediate layers (see Section 3.2).
Appendix B Small-Scale Ablation Studies
Before the large-scale training using the entire 180k hours of audio data, we conducted preliminary experiments on MuST-C v2 En-De (Cattoni et al., 2021) to investigate the effect of the CNN downsampling rate and the choice of the task for intermediate CTC layers. Specifically, we train 24-layer E-Branchformer-CTC models on the combined ASR and ST data from MuST-C v2 En-De. The input is always English audio, but the output can be the English ASR transcript or its German translation depending on the task specifier (see Figure 2).
@@ B.1 Effect of downsampling strategies
Table 14 compares different downsampling strategies while the other configurations are kept the same. The attention is implemented with FlashAttention (Dao et al., 2022). Self-conditioned CTC is applied at three intermediate layers: 6, 12, and 18. The first two CTC layers always perform ASR, while the others are task-dependent. The results show that using 8x downsampling in the CNN module leads to a slight degradation on WER and BLEU but reduces the GPU memory usage by half. We thus decide to employ 8x downsampling in our large-scale OWSM-CTC, enabling a doubled batch size per GPU. As mentioned in Appendix A.2, with this strategy, we observe a similar training speed compared to the encoder-decoder OWSM model.
@@ B.2 Choice of the CTC task
As discussed in Section 3.2, the intermediate CTC layers can be configured to perform a specific task like ASR or multiple tasks depending on the input task token. Table 15 compares different choices at a small scale using MuST-C v2 En-De. If all CTC layers are task-dependent (i.e., multi-tasking), the model cannot converge when trained from scratch. As more layers are used for ASR only, the ASR WER improves, but the ST BLEU decreases slightly. A good trade-off is to use the first half for ASR only and the second half for multi-tasking. Therefore, in our large-scale OWSM-CTC with 27 layers, we configure the 6th, 12th, and 15th layers to perform ASR only and the other two CTC layers (i.e., 21st and 27th layers) to be multi-tasking. This design also mimics the conventional cascaded system for ST.
Appendix C More Results of ST
Section 4.4 shows the BLEU scores using true case with punctuation. In this section, Table 16 and Table 17 present BLEU in lowercase without punctuation, which is consistent with the setup in prior work (Peng et al., 2024). The findings are very consistent with those in Section 4.4. Our OWSM-CTC achieves higher BLEU scores with faster inference speeds than the encoder-decoder OWSM v3.1 in general.
Appendix D Effect of text prompt
@@
Table 18 presents an example from TEDLIUM, where the text prompt changes the output style. When there is no prompt, the ASR output of OWSM-CTC is in true case with punctuation, and the apostrophes are combined with the previous words. However, when the previous sentence is used as a prompt, the style of the ASR hypothesis becomes more similar to that of the prompt. Specifically, the text is now in lowercase without punctuation marks, and the apostrophes are separate from previous words. This style is closer to the groundtruth transcript.
@@
Although the above example looks promising for biasing the model’s output toward certain directions, we note that this is not guaranteed to work in a zero-shot manner. We have also tried a few examples for zero-shot contextual biasing, where we provide a few biasing words in the prompt (e.g., person names), but we find that the model may not generate the correct word in many cases. This is mainly because the model is not really trained to perform this type of task - we just provide the previous sentence (according to some probability) as the prompt during training, which might not be useful at all; thus, the non-autoregressive model can simply ignore it in most cases. A more practical way to utilize this feature is to fine-tune our pre-trained model using some carefully designed data for contextual biasing. We will explore this in the future.
Appendix E Robustness